OpenStudy (anonymous):

Medal+Fan if answered with explanation on how to do this thanks! Please! Help me!

OpenStudy (anonymous):

3x+8y=12 2x+2y=3 1: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. 2: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

@ml1042903 could you please help me? my friend said you were good and very helpful

OpenStudy (anonymous):

hold up it might take a sec.

OpenStudy (anonymous):

thats okay.

OpenStudy (anonymous):

Thanks..

OpenStudy (anonymous):

@DE3100

OpenStudy (the_fizicx99):

This would be very easy if the drawing tool was working D: even though I'm a lousy artist

OpenStudy (the_fizicx99):

I can do it step by step but it'll look messy lol Cx

OpenStudy (anonymous):

Its fine.

OpenStudy (wolfe8):

So pick one equation and multiply it by a scalar. Can you do that?

OpenStudy (the_fizicx99):

I'll take the first equation, I'll multiply every term by -2 to make the x's the same. 3x(-2)+8y(-2)=12(-2) The new equation is -6x -16y = -24 For the other I'll multiply it by 3, 2x(3)+2y(3)=3(3) 6x +6y = 9 _________________________________________________ New equation: {-6x -16y = -24 { 6x +6y = 9 See that the x's cancel out? Add them up, -10y = -15 -10y = -15 /-10 /-10 y = 1.5 Plug in y (1.5) to solve for x, 6x +6(1.5) = 9 6x + 9 = 9 -9 -9 6x = 0 /6 /6 x = 0 So check my plugging in, 6(0) + 6(1.5) = 9 0 + 9 = 9 9 = 9 <-- this is true Now try this on the original equation, your pair is (0,1.5) 3(0)+8(1.5)=12 0 + 12 = 12 12 = 12 <-- true

OpenStudy (the_fizicx99):

By plugging in (0,1.5) into one of the equation it should give you a true statement :3

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (the_fizicx99):

You welcome :*

OpenStudy (anonymous):

Yea thanks @tHe_FiZiCx99 she really needed the help :)

OpenStudy (the_fizicx99):

>.< k